# RDT-1B: a Diffusion Foundation Model for Bimanual Manipulation

### v1

@@ 1 Introduction
However, it is highly non-trivial to develop a bimanual manipulation foundation model. One main reason is that the accessible data for a specific dual-arm robot is significantly scarce (Sharma et al., 2018; Collaboration et al., 2023) due to high hardware costs, undermining the data-intensive requirements of training foundational models. Inspired by recent attempts in unimanual manipulation (Brohan et al., 2023; Kim et al., 2024), we seek to first pre-train on extensive multi-robot datasets and then fine-tune on the small dataset collected on the target dual-arm robot. This can help us to scale the data size up to three orders of magnitude, having the potential to learn transferrable physics knowledge from datasets of other robots. Nevertheless, there are two key technical challenges. First, a generalizable foundation model requires a highly capable architecture in terms of both expressiveness and scalabilty. The dimension of the action space in bimanual manipulation is twice that in unimanual manipulation, bringing a higher degree of multi-modality in the distribution of feasible actions (Li, 2006; Jia et al., 2024), as illustrated in Fig. 2(b). Accordingly, the model must be expressive enough to capture the multi-modality in action distributions. Previous methods (Zhao et al., 2023; Brohan et al., 2023; Kim et al., 2024) typically fail to meet this standard, leading to unsatisfactory performance. Besides, the architecture needs to effectively process inputs from different modalities, including text, images, and actions. It must be scalable to stably train on large-scale robotic data. Second, date heterogeneity, which is caused by variations in the physical structure and the action space definition across different robots, can lead to negative transfer and impeding policy generalization during training on multi-robot data (Pan & Yang, 2009). Existing approaches either discard robots with differing action spaces or retain only the parts of the data whose structure is constant across the robot, at the cost of losing valuable data (Brohan et al., 2023; Ghosh et al., 2023; Shah et al., 2023a).
@@
In this paper, we introduce the Robotics Diffusion Transformer (RDT), the largest bimanual manipulation foundation model with strong generalizability. RDT employs Diffusion Transformers (DiTs) as its scalable backbone network (Xie et al., 2020), with special designs for language-conditioned bimanual manipulation with vision. For expressiveness, RDT excels in capturing the full modalities of bimanual actions from massive data by leveraging diffusion models’ capacity for modeling complex distributions (Sohn et al., 2015; Ho et al., 2020). For scalability, we harness the Transformer backbone and carefully design the multi-modal encoding to eliminate the heterogeneity of various modalities. To characterize the nonlinear dynamics (de Wit et al., 2012), high-frequency changes (Ghosh et al., 2023), and unstable numerical range inherent in robotic data, significantly differing from images and videos with temporal and spatial continuity (Chen et al., 2019; Liang et al., 2022), we make important modifications to the original DiT structure, including MLP decoding, improved normalization, and alternating condition injecting (See Fig. 4 for their importance). To further enable training RDT on heterogeneous data, we propose the Physically Interpretable Unified Action Space, a unified action format for various robots with gripper arms. This innovative format mitigates potential conflicts between different robots while retaining the physical meanings of the original actions, which can promote the model to learn generalizable physical knowledge across diverse robotic datasets.
@@ 2 Related Work
Learning-based Bimanual Manipulation. One substantial challenge in learning a bimanual manipulation policy is the high dimensionality of the action space, which exacerbates the data scarsity (Zollner et al., 2004; Smith et al., 2012; Lioutikov et al., 2016; Stepputtis et al., 2022) and the multi-modal behavior (Colomé & Torras, 2018; 2020; Figueroa & Billard, 2017; Sharma et al., 2018; Xie et al., 2020; Franzese et al., 2023). Some works have developed more cost-effective interfaces for data collection (Zhao et al., 2023; Aldaco et al., 2024), but they are limited to specific hardware configurations and still insufficient to bridge the data gap for a generalizable policy. Others attempt to reduce data requirements by introducing inductive biases, such as distinguishing two arms for stabilization and functionality (Grannen et al., 2023b), parameterizing movement primitives (Batinica et al., 2017; Amadio et al., 2019; Chitnis et al., 2020; Franzese et al., 2023), or using voxel representations (Grotz et al., 2024; Liu et al., 2024). These methods use strong priors or simplified modeling, which successfully reduce the action space, but at the cost of a reduced scope of application and failing to express the multi-modality of bimanual behaviors (Pearce et al., 2023).
Foundation Models for Robotics. Foundation models have shown immense promise in enabling generalizable behaviors by training multi-task “generalist” models (Brohan et al., 2022; 2023; Ghosh et al., 2023; Kim et al., 2024) on large multi-task robot datasets (Collaboration et al., 2023; Brohan et al., 2022; Fang et al., 2023). Most studies adapt large vision-language models to directly predict action (Brohan et al., 2022; Driess et al., 2023; Brohan et al., 2023; Collaboration et al., 2023; Kim et al., 2024). While demonstrating generalization to new objects and tasks, they face issues with quantization errors and uncoordinated behaviors (Pearce et al., 2023) when applied to bimanual manipulation, largely due to their discretization of action spaces. To enhance precision, diffusion models have been used for continuous control (Ho et al., 2020; Chi et al., 2023; Pearce et al., 2023; Ghosh et al., 2023). Ghosh et al. (2023) pre-train a Transformer-based diffusion policy on a subset of Open X-Embodiment (Collaboration et al., 2023) dataset ($25$ datasets), with up to $93$M parameters.
@@ 3 Problem Formulation and Challenges
We consider the concrete task of language-conditioned bimanual manipulation with vision, which is fundamental in robotics and has great value in real-world scenarios such as household (Stepputtis et al., 2020; Brohan et al., 2022; Zhao et al., 2023). Formally, given a language instruction $\ell$, the policy is presented with an observation ${\bm{o}}_{t}$ at time $t\in\mathbb{N}^{+}$; and then it produces an action ${\bm{a}}_{t}$ to control the two robot arms to achieve the goal specified by $\ell$. The observation is represented as a triple ${\bm{o}}_{t}:=({\bm{X}}_{t-T_{\text{img}}+1:t+1},{\bm{z}}_{t},c)$, where ${\bm{X}}_{t-T_{\text{img}}+1:t+1}:=({\bm{X}}_{t-T_{\text{img}}+1},\dots,{\bm{X}}_{t})$ is the RGB observation history of size $T_{\text{img}}$, ${\bm{z}}_{t}$ is the low-dimensional proprioception of the robot, and $c$ is the control frequency. The action ${\bm{a}}_{t}$ is usually a subset of the desired proprioception ${\bm{z}}_{t+1}$ [^1] E.g., ${\bm{z}}_{t}$ may include the gripper position at time $t$, and ${\bm{a}}_{t}$ can be the target gripper position at step $t+1$..
A specific task in bimanual manipulation typically consists of multiple elements: a skill (e.g., verbs like “pick” or “wipe”), an object (e.g., nouns like “bottle” or “table”), a scene (i.e. the environment in which the task takes place), and a modality describing how the skill is performed (e.g., adverbials like “pick the bottle with the left hand”). Given a new task, a practical policy is required to generalize to unseen elements that have not appeared in training data. This is challenging for previous rule-based methods as well as learning-based methods with small models/data, as discussed in Section 2.
@@ 4.1 RDT Model
Due to multi-modality, given the language instruction $\ell$ and observation ${\bm{o}}_{t}$, there may be many possible actions ${\bm{a}}_{t}$ to proceed with the task. The policy will learn the “average” of action modes if we model it as a deterministic mapping $(\ell,{\bm{o}}_{t})\mapsto{\bm{a}}_{t}$ and regress the tuples of $(\ell,{\bm{o}}_{t},{\bm{a}}_{t})$ in the training data. This may result in out-of-distribution actions, such as the arithmetic mean of multiple modes, which can be completely infeasible (Pearce et al., 2023). Instead, we choose to model the continuous conditional distribution $p({\bm{a}}_{t}|\ell,{\bm{o}}_{t})$. As discussed in Section 2, among various approaches, diffusion models excel in both expressiveness and sampling quality, but can be slow to sample a high-dimensional data (e.g., images). Luckily, for our settings, the drawback is minor since that ${\bm{a}}_{t}$ has a much lower dimension than images, which requires only minimal sampling overhead. This has made diffusion models an ideal choice for policy as in Chi et al. (2023).
@@
In addition to structure, heterogeneity features the different amounts of information contained in different inputs. First, data in different modalities contain different amounts of information. For example, images usually contain more information than text, and after encoding, they produce more tokens. Second, different inputs of the same modality may hold very different amounts of information. For example, the exterior camera of a robot has a more global view and contains richer information than the wrist cameras, as shown in the upper right of Fig. 3. In this case, the model may learn a shortcut: only focusing on the exterior view and ignoring the wrist views, thereby losing the ability to perceive depth. To tackle this issue, we randomly and independently mask each multi-modal input with a certain probability during encoding to prevent the model from over-relying on a specific input.
@@
  • Alternating Condition Injection (ACI). In our model, image and language inputs serve as conditions, which are high-dimensional and variable in length, contrasting with the class label conditions in traditional DiTs (Xie et al., 2020). These informative conditions are challenging to compress into a single token, making the original adaptive layer norm approach unsuitable. Therefore, we employ cross-attention to accommodate conditions of varying lengths avoiding the information loss in further compression. Besides, we further analyze that, given that image tokens are usually much more than text tokens, simultaneous injection of both modalities tends to overshadow text-related information, thus impairing the capability of the instruction following (see Fig. 4 for quantitative results). To mitigate this issue, we strategically alternate between injecting image and text tokens in successive layers’ cross-attention rather than injecting both in every layer.
@@ 4.2 Data
To enable training on heterogeneous multi-robot data, we need a unified action space shared among various robots, which can provide a unified format for multi-robot actions. The mapping from the original action space of a robot to the unified action space should be physically interpretable, each of whose dimensions should have a clear physical meaning. This can encourage the model to learn shared physical laws from different robot data, thereby improving the efficiency of learning from data of different robots (Shah et al., 2023a).
@@
Under this unified space, we are able to pre-train RDT on data from almost all modern robots with gripper arms and greatly expand the data scale toward the requirement for a foundation model. Specifically, our collection of pre-training datasets includes $46$ datasets of various robots, with a total size of $1$M+ trajectories and $21$TB. More details and preprocessing are deferred to App. D.
@@ 6 Conclusion
To reproduce our pre-training and fine-tuning processes, we have provided the code in the supplementary materials. We also include instructions for downloading the dataset, how to use the training code, and a guide for deploying on a real machine in the README file. Once the paper is accepted, we will fully open-source all our code, model weights, and fine-tuning datasets.
@@ Appendix H Experiment Details
The general performance in Fig. 1 of each method is calculated in three steps. Firstly, we calculate the success rate of a method in each task. We take an average of the total success rate and any additional requirement, i.e., the average of the values in the Total column and all columns to its right in Table 3. For example, in the Pouer Water-L-1/3, we take the average of Total, Correct Hand, and Correct Amount. Secondly, we calculate the success rate of each dimension of Unseen Object, Unseen Scene, Instruction Following, Few-Shot Learning, and Dexterity by averaging all the tasks in this dimension (see Table 1 for the correspondence). Lastly, we average the success rates of all the dimensions to obtain the overall result.

### v2

@@ 1 Introduction
Developing bimanual manipulation foundation models confronts the dual challenges of data scarcity and architectural limitations. The prohibitive costs of dual-arm systems create severe data scarcity (Sharma et al., 2018; Collaboration et al., 2023), fundamentally conflicting with the data-hungry nature of foundation models. Inspired by recent attempts in unimanual manipulation (Brohan et al., 2023; Kim et al., 2024), we mitigate this through cross-robot pretraining: leveraging multi-robot datasets for pre-training followed by target-robot fine-tuning, amplifying data volume by three orders of magnitude to extract transferable physical priors. However, two interconnected technical barriers emerge. First, the doubled action space induces multi-modal action distributions (Li, 2006; Jia et al., 2024) (see Fig. 2(b) for an illustrative example) that demand expressiveness capability beyond current methods (Zhao et al., 2023; Brohan et al., 2023; Kim et al., 2024), while simultaneously requiring scalabilty for stable large-scale training on multimodal data (text, vision, actions). Beyond architectural constraints, physical and action space variations across robots introduce data heterogeneity that risks negative transfer (Pan & Yang, 2009). Existing solutions either discard robots with structural inconsistencies or retain only cross-robot invariant features (Brohan et al., 2023; Ghosh et al., 2023; Shah et al., 2023a), sacrificing valuable data diversity essential for generalization.
@@
In this paper, we introduce the Robotics Diffusion Transformer (RDT), the largest bimanual manipulation foundation model with strong generalizability. RDT employs diffusion transformer (DiT) as its scalable backbone (Peebles & Xie, 2023), with special designs for language-conditioned bimanual manipulation. For expressiveness, RDT excels in capturing the full modalities of bimanual actions from massive data by using the capacity of diffusion models to represent complex distributions (Sohn et al., 2015; Ho et al., 2020). For scalability, we harness the Transformer backbone and carefully design the multi-modal encoding to eliminate the heterogeneity of various modalities. Moreover, robotic data is differed significantly from images and videos with temporal and spatial continuity (Chen et al., 2019; Liang et al., 2022). To characterize its inherent nonlinear dynamics (de Wit et al., 2012), high-frequency changes (Ghosh et al., 2023), and the unstable numerical range, we make important modifications to the original DiT structure, including MLP decoding, improved normalization, and alternate injection of conditions (see Fig. 4 for their importance). To further enable training RDT on heterogeneous data, we propose the Physically Interpretable Unified Action Space, a unified action format for various robots with gripper arms. This innovative format mitigates potential conflicts between different robots while retaining the physical meanings of the original actions, which can promote the model to learn generalizable physical knowledge across diverse robotic datasets.
@@ 2 Related Work
Learning-based Bimanual Manipulation. One substantial challenge in learning a bimanual manipulation policy is the high dimensionality of the action space, which exacerbates the data scarcity (Zollner et al., 2004; Smith et al., 2012; Lioutikov et al., 2016; Stepputtis et al., 2022) and the multi-modal behavior (Colomé & Torras, 2018; 2020; Figueroa & Billard, 2017; Sharma et al., 2018; Xie et al., 2020; Franzese et al., 2023). Some works have developed more cost-effective interfaces for data collection (Zhao et al., 2023; Aldaco et al., 2024), but they are limited to specific hardware configurations and still insufficient to bridge the data gap for a generalizable policy. Others attempt to reduce data requirements by introducing inductive biases, such as distinguishing two arms for stabilization and functionality (Grannen et al., 2023b), parameterizing movement primitives (Batinica et al., 2017; Amadio et al., 2019; Chitnis et al., 2020; Franzese et al., 2023), or using voxel representations (Grotz et al., 2024; Liu et al., 2024). These methods use strong priors or simplified modeling, which successfully reduce the action space, but at the cost of a reduced scope of application and inability to express the multi-modality of bimanual behaviors (Pearce et al., 2023).
Foundation Models for Robotics. Foundation models have shown immense promise in enabling generalizable behaviors by training multi-task “generalist” models (Brohan et al., 2022; 2023; Ghosh et al., 2023; Kim et al., 2024) on large multi-task robot datasets (Collaboration et al., 2023; Brohan et al., 2022; Fang et al., 2023). Most studies adapt large vision-language models to directly predict action (Brohan et al., 2022; Driess et al., 2023; Brohan et al., 2023; Collaboration et al., 2023; Kim et al., 2024). While demonstrating generalization to new objects and tasks, they face issues with quantization errors and uncoordinated behaviors (Pearce et al., 2023) when applied to bimanual manipulation. It’s largely due to their discretization of action spaces. To enhance precision, diffusion models have been used for continuous control (Ho et al., 2020; Chi et al., 2023; Pearce et al., 2023; Ghosh et al., 2023). Ghosh et al. (2023) pre-train a Transformer-based diffusion policy on a subset of Open X-Embodiment (Collaboration et al., 2023) dataset ($25$ datasets), with up to $93$M parameters.
@@ 3 Problem Formulation and Challenges
We consider the concrete task of language-conditioned bimanual manipulation with vision, which is fundamental in robotics and has great value in real-world scenarios such as household (Stepputtis et al., 2020; Brohan et al., 2022; Zhao et al., 2023). Formally, given a language instruction $\ell$, the policy is presented with an observation ${\bm{o}}_{t}$ at time $t\in\mathbb{N}^{+}$; and then it produces an action ${\bm{a}}_{t}$ to control two robot arms to achieve the goal specified by $\ell$. The observation is represented as a triple ${\bm{o}}_{t}:=({\bm{X}}_{t-T_{\text{img}}+1:t+1},{\bm{z}}_{t},c)$, where ${\bm{X}}_{t-T_{\text{img}}+1:t+1}:=({\bm{X}}_{t-T_{\text{img}}+1},\dots,{\bm{X}}_{t})$ is the RGB observation history of size $T_{\text{img}}$, ${\bm{z}}_{t}$ is the low-dimensional proprioception of the robot, and $c$ is the control frequency. The action ${\bm{a}}_{t}$ is usually a subset of the desired proprioception ${\bm{z}}_{t+1}$ [^1] E.g., ${\bm{z}}_{t}$ may include the gripper position at time $t$, and ${\bm{a}}_{t}$ can be the target gripper position at step $t+1$..
A specific task in bimanual manipulation typically consists of multiple elements: a skill (e.g., verbs like “pick”, “wipe”, or “open”), an object (e.g., nouns like “bottle”, “table”, or “door”), a scene (i.e. the environment in which the task takes place), and a modality describing how the skill is performed (e.g., adverbials like “pick the bottle with the left hand”). When encountering a new task, a practical policy is required to generalize to unseen222unseen means that a certain element has not appeared in the training data. elements in the task, which is particularly challenging for previous rule-based methods (Mirrazavi Salehian et al., 2017; Rakita et al., 2019; Grannen et al., 2023a) as well as learning-based methods that are limited to either small models and data or simple tasks, as discussed in Sec. 2.
@@ 4.1 RDT Model
Due to multi-modality, given the language instruction $\ell$ and observation ${\bm{o}}_{t}$, there may be many possible actions ${\bm{a}}_{t}$ to proceed with the task. The policy will learn the “average” of action modes if we model it as a deterministic mapping $(\ell,{\bm{o}}_{t})\mapsto{\bm{a}}_{t}$ and regress the tuples of $(\ell,{\bm{o}}_{t},{\bm{a}}_{t})$ in the training data. This may result in out-of-distribution actions, such as the arithmetic mean of multiple modes, which can be completely infeasible (Pearce et al., 2023). Instead, we choose to model the continuous conditional distribution $p({\bm{a}}_{t}|\ell,{\bm{o}}_{t})$. As discussed in Sec. 2, among various approaches, diffusion models excel in both expressiveness and sampling quality, but can be slow to sample high-dimensional data (e.g., images). Luckily, for our settings, the drawback is minor since that ${\bm{a}}_{t}$ has a much lower dimension than images, which requires only minimal sampling overhead. This has made diffusion models an ideal choice for policy as in Chi et al. (2023).
@@
Besides, heterogeneity also manifests in both cross-modal and intra-modal information density. First, modalities differ inherently in information capacity (e.g., visual inputs typically yield more tokens than text). Second, substantial variance exists within modalities, as seen in robotic perception where exterior cameras capture broader scenes versus the limited views from wrist cameras (Fig. 3). This discrepancy risks shortcut learning: focusing on exterior views while neglecting details from wrist cameras. To promote balanced multimodal integration, we implement stochastic independent masking across modalities during encoding, preventing overreliance on specific inputs.
@@
  • Alternating Condition Injection (ACI). In our model, image and language inputs serve as conditions, which are high-dimensional and variable in length, contrasting with the class label conditions in traditional DiTs (Peebles & Xie, 2023). These informative conditions are challenging to compress into a single token, making the original adaptive layer norm approach unsuitable. Therefore, we employ cross-attention to accommodate conditions of varying lengths avoiding the information loss in further compression. Besides, we further analyze that, given that image tokens are usually much more than text tokens, simultaneous injection of both modalities tends to overshadow text-related information, thus impairing the capability of the instruction following (see Fig. 4 for quantitative results). To mitigate this issue, we strategically alternate between injecting image and text tokens in successive layers’ cross-attention rather than injecting both in every layer.
@@ 4.2 Data
To enable training on heterogeneous multi-robot data, we need a unified action space shared among various robots to provide a unified format for multi-robot actions. The mapping from the original action space of a robot to the unified action space should be physically interpretable, and each dimension of the space should have a clear physical meaning. This can encourage the model to learn shared physical laws from different robot data, thereby improving the efficiency of learning from data of different robots (Shah et al., 2023a).
@@
With this unified space, we are able to pre-train RDT on data from almost all modern robots with gripper arms, and greatly expand the data scale towards the requirement for a foundation model. Specifically, our collection of pre-training datasets includes $46$ datasets of various robots, with a total size of $1$M+ trajectories and $21$TB. More details and preprocessing are deferred to App. D.
@@ 6 Conclusion
Acknowledgments
This work was supported by NSFC Projects (Nos. 92248303, 92370124, 62350080, 62276149, 62061136001, 92270001), BNRist (BNR2022RC01006), Tsinghua Institute for Guo Qiang, and the High Performance Computing Center, Tsinghua University. J. Zhu was also supported by the XPlorer Prize.
@@ Acknowledgments
To reproduce our pre-training and fine-tuning processes, we have provided the code in the repository. We also include instructions for downloading the dataset, how to use the training code, and a guide for deploying on a real machine in the README file. We have fully open-sourced all our code, model weights, and fine-tuning datasets. We refer to the project page for more information.
@@ Appendix H Experiment Details
The general performance in Fig. 1 of each method is calculated in three steps. Firstly, we calculate the success rate of a method in each task. We take an average of the total success rate and any additional requirement, i.e., the average of the values in the Total column and all columns to its right in Table 3. For example, in the Pour Water-L-1/3, we take the average of Total, Correct Hand, and Correct Amount. Secondly, we calculate the success rate of each dimension of Unseen Object, Unseen Scene, Instruction Following, Few-Shot Learning, and Dexterity by averaging all the tasks in this dimension (see Table 1 for the correspondence). Lastly, we average the success rates of all the dimensions to obtain the overall result.